# Landau and leading singularities in arbitrary space-time dimensions

### v2

@@ I Introduction
  • On top of the studies carried out for one-loop Feynman integrals, we employed the multivariate theory of residues given by Leray Leray_1959 in the calculation of leading singularities of two-loop planar and non-planar three-point functions with massive external and massless internal momenta. As a by-product of this direct calculation, we elaborated on the loop-by-loop approach and providing the $L$-loop ladder diagrams for three and four-point functions.
@@ II.2 One-loop decomposition
$=$ $\begin{vmatrix}p_{1}^{1}&\ldots&p_{D_{\parallel}}^{1}\\ p_{1}^{2}&\ldots&p_{D_{\parallel}}^{2}\\ \vdots&\ddots&\vdots\\ p_{1}^{D_{\parallel}}&\ldots&p_{D_{\parallel}}^{D_{\parallel}}\end{vmatrix}da_{1}\wedge\ldots\wedge da_{D_{\parallel}}\,.$ (18)
@@ II.3 Multi-loop decomposition
With the integration measure given by Eq. (24) in the expression of a generic multi-loop Feynman integral (1) at our disposal, an alert reader might notice a similarity with the Baikov representation. In effect, this derivation has been performed in Frellesvig:2017aai ; Primo:2017jtw , where the decomposition of space-time dimension, in parallel and perpendicular directions, was employed. Said differently, one obtains,
@@
Elaborating on this loop-momentum parametrization, we can proceed in the following to analyse Landau and leading singularities of different loop topologies. We shall recall that the physical information is contained in the parallel direction of the space-time dimension. The remaining integrations are simply cast in $\lambda_{ij}$. Therefore, we will mostly draw our attention to integrations in the parallel subspace.
@@ III Landau Singularities
Taking into account the independence of differential $d\ell$ the brackets in the last line of (III) give the well-known form of the Landau equations. In principle one can study singularities of subdiagrams by imposing $\alpha_{i}=0$ for some $i$, then Eq. (30a) in general should be stated as $D_{i}=0\text{ or }\alpha_{i}=0$. However, since we are interested in leading Landau singularities, all $\alpha$’s are non-zero.
@@
which should be understood that on-shell conditions and momentum conservation are taken into account. Then, the solution $\text{LanS}_{n}^{(1)}=0$ gives us the Landau singularity for one-loop diagrams. In this respect, Eq. (32) is given by the Gram determinant of internal momenta which can vanish identically if the number of internal momenta is larger than the dimension of the space. We will study the dimensional dependence of Landau singularities for one-loop diagrams.
@@ IV Leading Singularities
Additionally, because of the pattern displayed by explicit calculations of leading singularities in the various examples, we conjecture and prove the structure of these singularities in particular space-time dimensions – a connection between Landau and leading singularities. Similar results to those presented in this section can already be found in Abreu:2017ptx . However, whereas authors of the mentioned paper focused on dimensional regularisation and limited the analysis to a class of integrals where the integer dimension of the space-time is always even, omitting cases like the one-loop scalar triangle in the three space-time dimensions, we study the integrals in arbitrary space-time dimensions.
@@ IV.1 One-loop leading singularities
Feynman Integrals (IV.1) were obtained with the aid of relations (C), summarised in Appendix C. Also, to make explicit the parallel component of propagators, we define, $D_{i,\parallel}\equiv D_{i}-\lambda_{11}$, that is at most a quadratic polynomial in $a_{1}$ and absent of $\lambda_{11}$ (see Eq. (23)). To simplify the notation, here and in the following, we use the shorthand notation $D_{ij}=D_{i}-D_{j}$. The values of $r^{\pm}_{i}$, for $i=1,2$, are,
@@
Let us, however, remark that the integrands (IV.1), expressed as products of $d\log$ forms, are often referred to as integrands in $d\log$ representation and are by no means unique (see e.g. Henn:2021aco ). One can find alternative and more compact representations in terms of single $d\log$ integrands, by accounting for an educated change of variables. For instance,
@@
±$\pm\frac{1}{8\sqrt{-\text{LanS}_{3}^{\left(1\right)}}}\int_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathcal{K}}}d\log\frac{D_{1}}{D_{\pm}}\land d\log\frac{D_{2}}{D_{\pm}}\land d\log\frac{D_{3}}{D_{\pm}}\,,$ (63a)
@@
$\frac{1}{4\sqrt{\lambda_{\text{K}}\left(p_{1}^{2},p_{2}^{2},p_{3}^{2}\right)}}\int_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\mathcal{C}}}d\log D_{1}\land d\log D_{2}\land d\log D_{3}\,,$ (63b)
@@ IV.2 Connection between Landau and leading singularities
$\det(B)=$ $\begin{vmatrix}1&2\sum\limits_{j=1}^{n-1}p_{j}\cdot p_{1}a_{j}&\ldots&2\sum\limits_{j=1}^{n-1}p_{j}\cdot p_{n-1}a_{j}\\ 0&2p_{1}\cdot p_{1}&\ldots&2p_{1}\cdot p_{n-1}\\ \vdots&\dots&\vdots&\vdots\\ 0&2p_{n-2}\cdot p_{1}&\ldots&2p_{n-2}\cdot p_{n-1}\\ 0&2p_{n-1}\cdot p_{1}&\ldots&2p_{n-1}\cdot p_{n-1}\end{vmatrix}$ (85)
@@
$\det(B)$ $=(-1)^{2}2^{n-1}\begin{vmatrix}p_{1}\cdot p_{1}&\ldots&p_{1}\cdot p_{n-1}\\ \vdots&\dots&\vdots\\ p_{1}\cdot p_{n-2}&\ldots&p_{n-2}\cdot p_{n-1}\\ p_{1}\cdot p_{n-1}&\ldots&p_{n-1}\cdot p_{n-1}\end{vmatrix}$ (96)
@@ IV.3 Differential equations of one-loop Feynman integrals
In these illustrative examples, we showed that from the knowledge of master integrals in terms of one-loop scalar integrals, the construction of differential equations in the canonical form is automatic from the clear knowledge of leading singularities in specific space-time dimensions. Allowing, in this way, to focus more on the interesting properties of Feynman integrals given by their singularities.
@@ V.2 Loop-by-loop approach
From results (V.2), we observe that the presence of square roots in the denominator does not play any relevant role when computing the residue. We checked that these square roots can be rationalised and left the integrand in terms of only propagators. Therefore, we also obtain the four-dimensional leading singularity of a planar double box with massive external momenta and massless propagators (as depicted in Fig 3),
@@
With the results of Eqs. (V.2) and (V.2) at our disposal we can provide explicit expressions of leading singularities for ladder $L$-loop topologies, by exploiting the loop-by-loop approach. Our findings are summarised in the following theorems.
@@
[…]
@@
where by applying the results (V.2), we find,
@@
[…]
@@
Similar to theorem V.1, the proof is carried out by induction on $L$. For $L=2$, it was already computed in Eqs. (V.2). For $L=3$, let us start considering the three-loop Feynman integral,
@@
then, by using the results (V.2) for the two-loop ($L=2$) planar double box,
@@
After explicitly showing $L=2$ and $L=3$ cases, we can assume that our induction hypothesis is true for $L$ (with $L\geq 3$). Then, as similarly carried out in the previous theorem and making use of results (V.2), we find,
@@ VI Conclusions
Regarding leading singularities, we observed for one-loop scalar Feynman integrals that the number of integrations is lessened to the number of independent external momenta ($E$). This feature, obtained as a product of the general decomposition of the space-time dimension in terms of parallel and perpendicular directions, led us to perform an integrand analysis, finding a connection between leading and Landau singularities. In detail, we noted and proved that leading singularities in $D=E$ and $D=E+1$, respectively, correspond to the inverse of the square root of the leading Landau singularity of the first and second type. This connection was at first glance noted with the explicit calculation of leading and Landau singularities of up-to six-point one-loop scalar Feynman integrals. Then, to prove at all multiplicities, we use the multi-dimensional theory of residues provided by Leray.
@@
On top of the aforementioned one-loop analysis, we obtained, promising preliminary results for Feynman multi-loop integrals by applying developed methods. We considered particular families of integrals, namely $L$-loop three- and four-point ladder diagrams. Treating them as iterated integrals, we were able to apply Leray’s residues in the loop-by-loop approach to establish their leading singularities.

### v3

@@ I Introduction
  • On top of the studies carried out for one-loop Feynman integrals, we employed the multivariate theory of residues given by Leray Leray_1959 in the calculation of leading singularities of two-loop planar and non-planar three-point functions with massive external and massless internal momenta. As a by-product of this direct calculation, we elaborated on the loop-by-loop approach and providing the $L$-loop ladder diagrams for three and four-point functions. Continuing on the multi-loop application of the method we discuss the canonical differential equation for heavy-quark pair production through parton-parton annihilation in QCD at the two-loop level.
@@ II.2 One-loop decomposition
$=$ $\begin{vmatrix}p_{1}^{1}&\ldots&p_{D_{\parallel}}^{1}\\ p_{1}^{2}&\ldots&p_{D_{\parallel}}^{2}\\ \vdots&\ddots&\vdots\\ p_{1}^{D_{\parallel}}&\ldots&p_{D_{\parallel}}^{D_{\parallel}}\end{vmatrix}da_{1}\wedge\ldots\wedge da_{D_{\parallel}}\,.$
@@ II.3 Multi-loop decomposition
With the integration measure given by Eq. (19) in the expression of a generic multi-loop Feynman integral (1) at our disposal, an alert reader might notice a similarity with the Baikov representation. In effect, this derivation has been performed in Lee:2010wea ; Frellesvig:2017aai ; Primo:2017jtw . Within this framework, one obtains,
@@
Let us emphasize that the decomposition in terms of parallel and perpendicular components has been initially employed to derive Baikov representation of Feynman integrals to understand dimensional recurrence relations between Feynman integrals Lee:2010wea , alternatively to their parametric representation in terms of graph polynomials Tarasov:1996br ; Lee:2009dh . In the following, we make an extensive use of the representation (19) rather than Baikov representation because our main motivation lays in the mathematical study of the integration variables $a$ and $\lambda$.
Elaborating on this loop-momentum parametrization, we can proceed to analyse Landau and leading singularities of different loop topologies. We shall recall that the physical information is contained in the parallel direction of the space-time dimension. The remaining integrations are simply cast in $\lambda_{ij}$. Therefore, we will mostly draw our attention to integrations in the parallel subspace.
@@ III Landau Singularities
Taking into account the independence of differential $d\ell$ the brackets in the last line of (25) give the well-known form of the Landau equations. In principle one can study singularities of subdiagrams by imposing $\alpha_{i}=0$ for some $i$, then Eq. (25a) in general should be stated as $D_{i}=0\text{ or }\alpha_{i}=0$. However, since we are interested in leading Landau singularities, all $\alpha$’s are non-zero.
@@
which should be understood that on-shell conditions and momentum conservation are taken into account. Then, the solution $\text{LanS}_{n}^{(1)}=0$ gives us the Landau singularity for one-loop diagrams. This is a well established result know since the beginning of the analytic studies of Feynman integrals eden_1966 . In this respect, Eq. (27) is given by the Gram determinant of internal momenta which can vanish identically if the number of internal momenta is larger than the dimension of the space. We will study the dimensional dependence of Landau singularities for one-loop diagrams.
@@ IV Leading Singularities
Additionally, because of the pattern displayed by explicit calculations of leading singularities in the various examples, we conjecture and prove the structure of these singularities in particular space-time dimensions – a connection between Landau and leading singularities. Similar results to those presented in this section can already be found in Abreu:2017ptx which were obtained by working in the embedding formalism. Whilst authors of the mentioned paper focused on dimensional regularisation and present explicitly only analysis and results for even dimensions of the space-time, the odd dimensional cases can also by obtained by the method of Abreu:2017ptx by setting $\epsilon=-\frac{1}{2}$  [^4] The closely related notion of maximal cut integrals presented in  Abreu:2017ptx is valid in arbitrary $D$ dimensions. .
@@ IV.1 One-loop leading singularities
Feynman Integrals (54) were obtained with the aid of relations (153), summarised in Appendix C. Also, to make explicit the parallel component of propagators, we define, $D_{i,\parallel}\equiv D_{i}-\lambda_{11}$, that is at most a quadratic polynomial in $a_{1}$ and absent of $\lambda_{11}$ (see Eq. (18)). To simplify the notation, here and in the following, we use the shorthand notation $D_{ij}=D_{i}-D_{j}$. The values of $r^{\pm}_{i}$, for $i=1,2$, are,
@@
Let us, however, remark that the integrands (54), expressed as products of $d\log$ forms, are often referred to as integrands in $d\log$ representation and are by no means unique (see e.g. Henn:2021aco ). One can find alternative and more compact representations in terms of single $d\log$ integrands, by accounting for an educated change of variables. For instance,
@@
±$\pm\frac{1}{8\sqrt{-\text{LanS}_{3}^{\left(1\right)}}}\int_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{K}}}d\log\frac{D_{1}}{D_{\pm}}\land d\log\frac{D_{2}}{D_{\pm}}\land d\log\frac{D_{3}}{D_{\pm}}\,,$ (58a)
@@
$\frac{1}{4\sqrt{\lambda_{\text{K}}\left(p_{1}^{2},p_{2}^{2},p_{3}^{2}\right)}}\int_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathcal{C}}}d\log D_{1}\land d\log D_{2}\land d\log D_{3}\,,$ (58b)
@@ IV.2 Connection between Landau and leading singularities
$\det(B)=$ $\begin{vmatrix}1&2\sum\limits_{j=1}^{n-1}p_{j}\cdot p_{1}a_{j}&\ldots&2\sum\limits_{j=1}^{n-1}p_{j}\cdot p_{n-1}a_{j}\\ 0&2p_{1}\cdot p_{1}&\ldots&2p_{1}\cdot p_{n-1}\\ \vdots&\dots&\vdots&\vdots\\ 0&2p_{n-2}\cdot p_{1}&\ldots&2p_{n-2}\cdot p_{n-1}\\ 0&2p_{n-1}\cdot p_{1}&\ldots&2p_{n-1}\cdot p_{n-1}\end{vmatrix}$
@@
$\det(B)$ $=(-1)^{2}2^{n-1}\begin{vmatrix}p_{1}\cdot p_{1}&\ldots&p_{1}\cdot p_{n-1}\\ \vdots&\dots&\vdots\\ p_{1}\cdot p_{n-2}&\ldots&p_{n-2}\cdot p_{n-1}\\ p_{1}\cdot p_{n-1}&\ldots&p_{n-1}\cdot p_{n-1}\end{vmatrix}$
@@ IV.3 Differential equations of one-loop Feynman integrals
In these illustrative examples, we showed that from the knowledge of master integrals in terms of one-loop scalar integrals, the construction of differential equations in the canonical form is automatic from the clear knowledge of leading singularities in specific space-time dimensions.
The results of this section show that the construction of the canonical form of the differential equation for one-loop Feynman integrals is completely determined by the Landau analysis of master integrals. Allowing, in this way, to focus more on the interesting properties of Feynman integrals given by their singularities.
@@ V.2 Loop-by-loop approach
From results (98), we observe that the presence of square roots in the denominator does not play any relevant role when computing the residue. We checked that these square roots can be rationalised and left the integrand in terms of only propagators. Therefore, we also obtain the four-dimensional leading singularity of a planar double box with massive external momenta and massless propagators (as depicted in Fig 3),
@@
With the results of Eqs. (98) and (99) at our disposal we can provide explicit expressions of leading singularities for ladder $L$-loop topologies, by exploiting the loop-by-loop approach. Our findings are summarised in the following theorems.
@@
The leading singularity of the three-point $L$-loop ladder Feynman integral of Fig. 4 in four space-time dimensions with off-shell external momenta ($p_{i}^{2}\neq 0$ for $i=1,2,3$) and massless propagators is equal to $\left(\left(p_{3}^{2}\right)^{L}\sqrt{\lambda_{\text{K}}\left(1,\frac{p_{1}^{2}}{p_{3}^{2}},\frac{p_{2}^{2}}{p_{3}^{2}}\right)}\right)^{-1}$.
@@
where by applying the results (98), we find,
@@
The leading singularity of the four-point $L$-loop ladder Feynman integral of Fig. 4 in four space-time dimensions with off-shell external momenta ($p_{i}^{2}\neq 0$ for $i=1,2,3,4$) and massless propagators is equal to $\left(s^{L}t\sqrt{\lambda_{\text{K}}\left(1,\frac{p_{1}^{2}p_{3}^{2}}{st},\frac{p_{2}^{2}p_{4}^{2}}{st}\right)}\right)^{-1}$, with $s=(p_{1}+p_{2})^{2}$ and $t=(p_{2}+p_{3})^{2}$.
@@
Similar to theorem V.1, the proof is carried out by induction on $L$. For $L=2$, it was already computed in Eqs. (99). For $L=3$, let us start considering the three-loop Feynman integral,
@@
then, by using the results (99) for the two-loop ($L=2$) planar double box,
@@
After explicitly showing $L=2$ and $L=3$ cases, we can assume that our induction hypothesis is true for $L$ (with $L\geq 3$). Then, as similarly carried out in the previous theorem and making use of results (99), we find,
@@
V.3 Observations on differential equations in canonical form
In view of the loop-by-loop approach considered in Sec. V.2, obtained from leading and Landau singularities of Feynman integrals at one-loop, we can combine this approach with evaluation of integrals that are needed for the calculation of scattering amplitudes at high energies. Previously, we remarked that the knowledge of $d\log$ integrals in the top sector in interplay with the algorithm Dlapa:2020cwj can automatically lead us to construct a canonical basis of master integrals. In this section, we plan to avoid the use of the latter algorithm and profit, instead, from the study of multi-loop Feynman integrals with constant leading singularity in a given dimension. To complete the canonical basis, we use the knowledge of leading singularities in different dimensions, extending the discussion of Sec. IV.3.
Figure 5: One- and two-loop integral families. Thin (thick) lines correspond to massless (massive) propagators
To start, let us draw our attention to the integral families depicted in Fig. 5. These integrals appear in the calculation of one- and two-loop scattering amplitudes for heavy-quark pair production through parton-parton annihilation in QCD. We can then express these families, respectively, as,
$J_{4}^{\left(1\right),D}\left(1^{n_{1}},2^{n_{2}},3^{n_{3}},4^{n_{4}}\right)=$ $\int\frac{d^{D}\ell_{1}}{\imath\,\pi^{D/2}}\frac{1}{\left(\ell_{1}-p_{1}\right)^{2n_{1}}\ell_{1}^{2n_{2}}\left(\ell_{2}+p_{2}\right)^{2n_{3}}\left(\left(\ell_{2}+p_{23}\right)^{2}-m^{2}\right)^{n_{4}}}\,,$ (108)
$J_{4}^{\left(1\right),D}\left(1^{n_{1}},\ldots,7^{n_{7}};8^{n_{8}},9^{n_{9}}\right)=$ $\int\frac{d^{D}\ell_{1}}{\imath\,\pi^{D/2}}\frac{d^{D}\ell_{2}}{\imath\,\pi^{D/2}}\frac{1}{\left(\ell_{1}-p_{1}\right)^{2n_{1}}\ell_{1}^{2n_{2}}\left(\ell_{1}+p_{2}\right)^{2n_{3}}}$
×$\times\frac{\left(\left(\ell_{1}+p_{23}\right)^{2}-m_{2}^{2}\right)^{-n_{8}}\left(\ell_{2}^{2}\right)^{-n_{9}}}{\left(\ell_{2}-p_{1}\right)^{2n_{4}}\left(\ell_{2}+p_{2}\right)^{2n_{5}}\left(\left(\ell_{2}+p_{23}\right)^{2}-m^{2}\right)^{n_{6}}\left(\ell_{1}-\ell_{2}\right)^{2n_{7}}}\,,$ (109)
with $p_{1}^{2}=p_{2}^{2}=0$, $p_{3}^{2}=p_{4}^{2}=m^{2}$, the momentum conservation $p_{1}+p_{2}+p_{3}+p_{4}=0$, and the kinematic invariants, $s=(p_{1}+p_{2})^{2},t=(p_{2}+p_{4})^{2}$. We follow the convention of (1). These integral families are known to have, respectively, 5 and 21 master integrals. While at one-loop it is straightforward to obtain a canonical basis from the knowledge of leading and Landau singularities in different dimensions (see Sec. IV.3), for the two-loop case, instead, a loop-by-loop approach can be considered. We follow an alternative approach based solely on the study of $d\log$ integrals in a specific dimension. Without loss of generality, we work on $D=4-2\epsilon$.
As considered throughout this work, looking for $d\log$ integrals relies on the loop-momentum parametrisation of the integrand. Thus, by following the approach of Henn:2020lye , one can construct an ansatz of possible candidates of $d\log$ integrals to then solve for basis of integrals that only contain a constant leading singularity. This method has been successfully developed in the code DlogBasis by profiting from the spinor-helicity formalism for the parametrisation of the loop momenta.
Figure 6: Missing canonical master integrals.
Owing to the DlogBasis implementation, we compute candidates for the one- and two-loop integral families of Fig. 5. We realise that, after IBP reduction, there are missing integrals in the sectors displayed in Fig. 6. These integrals appear in subsectors and only contain insertions of bubble and tadpole loops. We consider these sub-loop integral in $D=2-2\epsilon$, whilst the other loop remains in $D=4-2\epsilon$. Let us look at this in detail.
For the integrals displayed in Fig. 6, we find the canonical integrals to admit the following form,
$J_{4}^{\left(1\right),D=4-2\epsilon}\left(4\right)$ $\to J_{4}^{\left(1\right),D=2-2\epsilon}\left(4\right)\,,$
$J_{4}^{\left(2\right),D=4-2\epsilon}\left(3,4,6,7\right)$ $\to r_{\lambda}\,J_{4}^{\left(2\right),D=4-2\epsilon}\left(3,4,6,7^{2}\right)\,,$
$J_{4}^{\left(2\right),D=4-2\epsilon}\left(1,3,6,7\right)$ $\to s\,J_{4}^{\left(2\right),D=2-2\epsilon}\left(1,3,6,7;8^{-1}\right)\,,$
$J_{4}^{\left(2\right),D=4-2\epsilon}\left(3,6,7\right)$ $\to m^{2}\,J_{4}^{\left(2\right),D=4-2\epsilon}\left(3,6,7;8^{-1}\right)\,,$
$J_{4}^{\left(2\right),D=4-2\epsilon}\left(3,4,7\right)$ $\to s\,J_{4}^{\left(2\right),D=2-2\epsilon}\left(3,4,7\right)$
$J_{4}^{\left(2\right),D=4-2\epsilon}\left(2,6,7\right)$ $\to t\,J_{4}^{\left(2\right),D=2-2\epsilon}\left(2,6,7;8^{-1}\right)$
$J_{4}^{\left(2\right),D=4-2\epsilon}\left(1,3,6\right)$ $\to s\,J_{4}^{\left(2\right),D=2-2\epsilon}\left(1,3,6\right)\,,$
with $r_{\lambda}=\sqrt{\lambda_{\text{K}}\left(s,m^{2},m^{2}\right)}=\sqrt{-s\left(-s+4m^{2}\right)}$. Integrals that only contain a bubble sub-loop are independently treated in $D=2-2\epsilon$ and then shifted to $D=4-2\epsilon$. Thus, introducing squared propagators. For the canonical integrals that contain an insertion of a numerator, we directly consider them in $D=2-2\epsilon$. With this additional set of master integrals, we recover the canonical differential equations.
@@ VI Conclusions
Regarding leading singularities, we observed for one-loop scalar Feynman integrals that the number of integrations is lessened to the number of independent external momenta ($E$). This feature, obtained as a product of the general decomposition of the space-time dimension in terms of parallel and perpendicular directions, led us to perform an integrand analysis, finding a connection between leading and Landau singularities. In detail, we noted and proved that leading singularities in $D=E$ and $D=E+1$, respectively, correspond to the inverse of the square root of the leading Landau singularity of the first and second type. At the beginning we observed this connection with the explicit calculation of leading and Landau singularities of up-to six-point one-loop scalar Feynman integrals. Then, to prove at all multiplicities, we use the multi-dimensional theory of residues provided by Leray.
@@
On top of the aforementioned one-loop analysis, we obtained, promising preliminary results for Feynman multi-loop integrals by applying developed methods. We considered particular families of integrals, namely $L$-loop three- and four-point ladder diagrams. Treating them as iterated integrals, we were able to apply Leray’s residues in the loop-by-loop approach to establish their leading singularities. Continuing on the multi-loop application of the method we discussed the canonical differential equation for heavy-quark pair production through parton-parton annihilation in QCD at the two-loop level.